OpenStudy (anonymous):

Which set of line segments with the given lengths can be formed into a right triangle? 24,30,35 12,18,30 18,24,30 18,24,35

OpenStudy (amistre64):

gonna have to pythag it

OpenStudy (lalaly):

the third

OpenStudy (amistre64):

18,24,30/3 6,8,10 .... 3,4,5 thats it :) good eye lalaly

OpenStudy (angela210793):

18^2+24^2=30^2 144+576=900 900=900 Lana's right :)